40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Mar 8 03:27:48 2015, Jenne, Update, LSC, Error signal blending for CARM/DARM transitions 
    Reply  Sun Mar 8 04:04:19 2015, Jenne, Update, LSC, Error signal blending for CARM/DARM transitions 
       Reply  Sun Mar 8 13:51:41 2015, rana, Update, LSC, Error signal blending for CARM/DARM transitions 
          Reply  Mon Mar 9 14:14:32 2015, Jenne, Update, LSC, Error signal blending for CARM/DARM transitions First-ever_RFonly_lock_7March2015.pngSecond-ever_RFonly_lock_7March2015.png
Message ID: 11119     Entry time: Sun Mar 8 03:27:48 2015     Reply to this: 11120
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Error signal blending for CARM/DARM transitions 

This elog will be about work that happened yesterday.  I will write a reply to this with work from this evening's success.


[Rana, Jenne]

Work started with the plan of trying ALS fool, using the new triggering scheme (elog 11114).

The PRMI was having a bit of trouble holding lock with REFL165, so we checked its demod phase.  On Monday (elog 11095) we rotated the REFL165 phase from -91 deg to -48 deg while in PRFPMI configuration (I think the -91 was from PRMI-only phase setting).  However, Friday night we saw that MICH was super noisy, especially when the CARM and DARM offsets were near zero.  Rana rotated REFL165's phase until the MICH noise seemed to get lower (by at least an order of magnitude in the control signal), while we were at zero offset everywhere. We were not driving and looking at any lines/peaks, just the overall spectra.  The final REFL165 demod phase is -80. 

We tried engaging the fool path with no success. 

First, Rana moved the low frequency boost in the MC filter bank from 20:1 to 0.3:0.03.  This gave the whole loop at least 20 or 30 degrees of phase at all frequencies below the design UGF (a few hundred Hz?  Don't quite remember).  To check this, we put in a "plant" filter, and turned on the locking filter (3:3000^2) and the low freq boost and the plant, and the phase never touched 180 at any low freq.  This is so that we can ramp on this filter bank's gain without having an unstable unity gain crossing anywhere.  Also, I added two +10dB filters to the first two filter modules, so that we could ramp on the gain at the input rather than the output.

Last night we were actuating CARM on MC2 and DARM on the ETMs, and the MC filter bank was set to actuate on MC2.  Even with super duper low gain in the MC filter bank, so that the control signal was much less than one (1) count, it would make CARM unhappy.  The CARM filter bank's output was doing +/- a hundred or more counts, so why a few tenths of a count mattered, we couldn't figure out.  We were using the power trigger for the MC filter bank, but not the zero-crossing trigger.  Since the fool tuning was checked while actuating on the ETMs, we wonder if maybe the tuning isn't valid for MC2 actuation?  Maybe there's enough of a difference between them that the fool needs to be re-tuned for MC2 actuation?  Fool had the complex pair of poles at 1Hz, the "comp1" filter to give phase lag, and a gain of 22. 

I think that at some point we even turned off the fool path, but left the MC path on with a little bit of gain, and the audible noise over the speakers didn't seem to change in character at all. Weird. 


We ended up leaving the fool path for another time, and started working on error signal blending at the CARM filter bank input.  This is pretty similar to Kiwamu's self-locking principle.

Our goal was to ramp up the gain of the RF error signal at low frequency, while letting ALS keep hold of things at higher frequencies. 

CARM and DARM sweeps from earlier seemed to indicate that the RF signals are valid without normalization above transmitted powers of 50 or so, so we thought we'd give those a whirl for this error signal blending.

From doing a CARM sweep through resonance, we guessed roughly that the REFL11 (non-normalized) slope was about a factor of 10,000 larger than the ALS slope.  We put a 1e-4 into the input matrix element REFL11I -> CARM_B.  For some reason, REFL11 seemed to be centered around -250 counts, so we put an offset of +0.025 ( = 250*1e-4) into the CARM_B filter module to compensate for this. 

Since we thought that a gain of 1 in the CARM_B filter bank would make it equal to ALS, we tried some lower gains to start with.  0.3 kicked it out of lock, so we ended up liking and using 0.15.  With this low gain on, we tried turning on a low frequency boost, 20:1, but that didn't do very much.  We turned that off, and instead turned on an integrator, 20:0, which totally made things better.  The transmitted arm power was staying higher more of the time.

From a DARM sweep, we thought that AS55Q (non-normalized) should also have an input matrix element of 1e-4 for DARM.  We gave DARM_B a gain of 0.1, which seemed good and not too high.  Again, trying the gentle boost didn't do much, so we went with the integrator. 

At this point, since both RF signals were being used as error signals with integrators, we declared that at least at DC we were on RF signals.  Hooray!! 

After this, we started increasing the CARM_B gain a little, and decreasing the CARM_A gain.  When Rana finally set the CARM_A element to zero, we lost lock.  We realized that this is because we didn't include a zero to compensate for the arm cavity pole, which the IR signal will see, but the ALS won't. 

We decided that the plan of attack would be to get back to where we were (DC error signals on RF), and try to start engaging the AO path. 

 

ELOG V3.1.3-